OpenStudy (anonymous):

4 over 5 x + 4 over 3 = 2x

OpenStudy (anonymous):

\[\frac{ 4 }{ 5x }+\frac{ 4 }{ 3 }=2x\]

OpenStudy (anonymous):

\(\Large \color{blue}{\frac{4}{5x}+\frac{4}{3}=2x}\\\Large \color{blue}{2x-\frac{4}{5x}=\frac{4}{3}}\\\Large \color{blue}{\frac{2x\times 5x-4}{5x}=\frac{4}{3}\\\Large \frac{10x^2-4}{5x}=\frac{4}{3}\\\Large 3\times (10x^2-4)=4\times 5x\\30x^2-12=20x\\\Large 30x^2-20x-12=0\\\Large 15x^2-10x-6=0}\) \(\Large \color{red}{\Rightarrow x=\frac{10\pm \sqrt{10^2+4\times 15\times 6}}{2\times15}=\frac{5\pm \sqrt{115}}{15}}\)